A student is riding their bike. They start by riding their bike east for 34 meters. They turn north and ride their bike 67 meters. They turn south and continue for 14 meters. Finally, they turn west and ride their bike 34 meters.

The displacement of the student is:

  Δ x=  unit

And the distance traveled is:

d =  unit
Answer #1

Ans : ∆x = 53 m , d = 149 m

WEE 14m nk 34 m 67m 34 m Displacment. - 4x = AB - BC. = 67 m- 14m => [ax=53m Ans. Distance travelled : d=OA+ AB+B C + CD = 34